# Towards an Agnostic Algorithm for Sampling Empirical Structure Models

### v1

@@
We present an algorithm to efficiently sample the full space of planetary interior density profiles. Our approach uses as few assumptions as possible to pursue an agnostic algorithm. The algorithm avoid the common Markov Chain Monte Carlo method for an optimisation-based gradient descent approach designed for computational efficiency. In this work, we use Uranus and Neptune as test cases and obtain empirical models that provide density and pressure profiles consistent with the observed physical properties (total mass, radius, and gravitational moments). We compare our findings to other work and find that while other studies are generally in line with our findings, they do not cover the entire space of solutions faithfully. Furthermore, we present guidance for modellers that construct Uranus or Neptune interior models with a fixed number of layers. We provide a statistical relation between the steepness classifying a density discontinuity and the resulting number of discontinuities to be expected. For example, if one classifies a discontinuity as a density gradient larger than 0.02 kg m-4, then most solutions should have at most one such discontinuity. Finally, we find that discontinuities, if present, are concentrated around a planetary normalised radius of 0.65 for Uranus and 0.7 for Neptune. Our algorithm to efficiently and faithfully investigate the full space of possible interior density profiles can be used to study all planetary objects with gravitational field data.
@@ 1 Introduction
Uranus and Neptune have always been key to understanding the formation and evolution history of the Solar System. As we continue to expand our investigation of our own planetary system, the study of exoplanets is gaining more attention. Although the characterization of exoplanets remains limited, it is now clear that in terms of mass and size, many exoplanets in the galaxy resemble Uranus and Neptune (Zhu & Dong, 2021). By revealing the internal structures of Uranus and Neptune, we can improve our knowledge of both these planetary objects in our own neighbourhood and the many observed intermediate-mass exoplanets orbiting distant stars.
To study the internal structure of a planet, we need to know its density profile. Lacking direct measurements, we need to infer this profile from other observational data. Density profiles can be inferred by varying all values constituting a profile until the result matches the available observational data (for example, total mass, radius, and the gravitational field). The number of such observational parameters is typically small, for Uranus and Neptune fewer than ten. In contrast, density profiles typically consist of orders of magnitude more data points. Inferring density profiles is therefore severely degenerate, many density profiles can fit the same observational data.
To address this problem, one can simplify the approach used to infer density profiles by relying on a multitude of assumptions. However, caution must be taken: Making unsound simplifying assumptions could inadvertently exclude all (or a significant portion of) the valid density profiles. It is therefore paramount to obtain a complete overview of all possible solutions before simplifications even considered.
@@
The current numerical method of choice for empirical structure modelling is often the Markov Chain Monte Carlo (MCMC) method (for example Marley et al., 1995; Movshovitz & Fortney, 2022). The MCMC approach starts by generating initial density profiles, and then randomly alters their parameters until they fit the observational data. This poses challenges due to the aforementioned high amount of data points that make up a density profile. One must therefore either tolerate a slow convergence of the MCMC method when handling countless parameters, or reduce the dimensionality of the parameter space to accelerate the search. Previous studies have taken the latter route by parametrising density profiles with polytropes, polynomials, or sigmoid functions. This reduces the number of parameters to an order of magnitude comparable to the number of observational parameters, but can introduce biases, precisely what we aim to avoid here.
We replace the common MCMC method with a new, optimisation-based approach, an algorithm to efficiently investigate the full space of possible interior density profiles that fit the known observational data of a planet. The assumptions made for the profiles are reduced to an absolute minimum: All we require is that there are no density plateaus or inversions and that the maximal density does not exceed the maximal density of any feasible material. Otherwise (with a resolution of $N=1024$) the $N$ density values are all allowed to change freely. Because we treat all $N$ density values as free parameters, the method avoids biases introduced by the parametrisations of previous MCMC studies. However, this alone does not ensure that the results are free from bias. We must further ensure that the generation of starting profiles is unbiased and — because we no longer use an MCMC algorithm — we must additionally ensure that the way we find the density profiles is unbiased. For the first point, if the generation of starting profiles favours certain types of density profiles, for instance, by initially under-representing profiles with sharp density discontinuities, an adjustment that increases the sampling of such profiles is required. For the second point, we need to verify that our optimisation approach does not overlook harder to find local solutions in favour of dominant global ones that are easily spotted.
Our paper is structured as follows. Section 2 provides an overview of the algorithm. Section 3 presents and interprets the results, Section 4 discusses the results, highlights caveats, and provides an outlook for future work, and Section 5 summarises the key findings of this study. Note that our methodology is based on the Theory of Figures (Zharkov & Trubitsyn, 1978). For the purposes of notation, we follow the elaboration presented in Appendix A of Morf et al. (2024). We employ the open-source code PyToF [^1] See https://doi.org/10.5281/zenodo.17592269 (Morf, 2025) to implement the Theory of Figures numerically.
@@ 2.1 Algorithm
Planets are flattened due to their rotation. Surfaces of both constant density and potential are not spherical shells defined by a single radius $r$, but rather spheroidal in shape. Spheroidal surfaces are described by a function $r_{l}(\vartheta)$ which depends on the polar angle $\vartheta$. The variable $l$ uniquely identifies a spheroid and represents its volumetric mean radius:
@@
where $R_{\text{ref}}$ is a normalisation constant, $m$ denotes the planetary mass, and $P_{2n}$ are the Legendre polynomials. The $J_{2n}$ are called gravitational moments and can be calculated with the Theory of Figures.
For the numerical implementation, density profiles consist of $N$ density values $\rho_{k}$ at evenly spaced level surfaces $l_{k}$. We restrict the maximum density to some threshold $\rho_{max}$ and enforce a minimum density increase per step $\Delta_{\text{min}}$. The density values are derived from N1N-1 parameters $p_{i}$ as
@@
The weights $w_{i}$ were added to normalise the parameter values and are defined as
@@
The scaling factor $\alpha$ was added to ensure the density profile mass $m_{calc}$ agrees with the planetary mass $m$. It is defined as
@@
where the calculated mass $m_{calc}$ was obtained by numerically integrating the density values over the level surfaces via
@@
Listing 1: The Optimisation Algorithm
@@
Given the number of epochs to run the algorithm for, we first create some suitable starting parameters. Then, for each epoch, we check whether the optimisation has converged to a stable solution. If not, we perform epochsize many optimisation steps. For each optimisation step, we calculate the gradient and move some amount along it according to the Adam gradient descent scheme (Kingma & Ba, 2017). Then, we project the new parameters to enforce our minimum density increase. Once convergence is detected, the final result, along with other data, is returned. Further details on the algorithm can be found in the Appendix A. Elaborations on the requirements of an unbiased generation method and optimisation method are given in Appendix C. The method’s limitations are discussed in Section 4.
Table 1: Physical values for Uranus and Neptune
@@ 2.2 Simulation setup
The simulation was performed on an AMD EPYC 7702 Hypervisor CPU with 96 cores over 120 hours. 64 cores were dedicated to Uranus and 32 towards Neptune. We have dedicated twice as many cores to Uranus in order to run a correlated ($\text{cov}(J_{2},J_{4})=0.9861\sigma_{J_{2}}\sigma_{J_{4}}$, see Table 1) and uncorrelated ($\text{cov}(J_{2},J_{4})=0$) simulation in parallel. All physical values for Uranus and Neptune are given in Table 1. The number of layers $N$ was set to 1024 and the epoch size was 25 steps. We chose a maximum of 120 epochs (3000 steps). The scalar factor $c$ introduced in Equation 10 was set to $100$. This factor ensures that the optimisation cost function does not neglect deviations in total mass in favour of optimising for gravitational field data. As a convergence criterion within the Theory of Figures algorithm PyToF, we employed $\left|J_{2n}^{i+1}-J_{2n}^{i}\right|\leq 10^{-10}J_{2n}^{i}$ to check if convergence has been achieved after the $i$-th iteration. The maximum acceptable density $\rho_{max}$ was set to 20000 kg m-3. The minimum acceptable density increase per unit length $\Delta_{\text{min}}$ was set to 1 $\Delta\ell^{-1}$ kg m-3, where $\Delta\ell=R/N$. Explicitly, this results in a value of $\Delta\ell=$ 24960 m for Uranus and $\Delta\ell=$ 24186 m for Neptune (rounded to the nearest meter). This yields $\Delta_{\text{min}}=$ 4.01 $\cdot$ 10-5 kg m-4 for Uranus and $\Delta_{\text{min}}=$ 4.13 $\cdot$ 10-5 kg m-4 for Neptune (rounded to two significant figures).
@@
Overall, 694465 (791901) runs were completed for the correlated (uncorrelated) Uranus data, and 731407 runs were completed for Neptune. For Uranus, 90.39% (90.48%) of the runs respected $\rho_{max}$, and 18.77% (18.56%) achieved gravitational moments $J_{2n}$ within the multivariate three standard deviation threshold, which is 20.77% (20.56%) of those that respected $\rho_{max}$. For Neptune, the same numbers are 89.22% and 88.83%, which is 99.57% of those that respected $\rho_{max}$. Our optimisation-based approach and the developed algorithm are highly efficient: The average time to complete a run was approximately 20 seconds.
@@
Note that throughout all subsequent results, the parameter space of possible density profiles for Neptune is larger in comparison to Uranus. This is because the measurements we have for the gravitational moments of Uranus are about one order of magnitude more precise than those for Neptune (Table 1). Consequently, Uranian runs had a $\sim$19% success rate, compared to $\sim$89% of the Neptunian runs.
@@
Figure 1: Gravitational moment distribution for successful Uranus (1(a)) and Neptune (1(b)) profiles. Both panels include a two-dimensional histogram where bins more frequently occupied by solutions are shown with relatively darker colour. Additionally, one-dimensional histograms display the weighted number of solutions along each axis. The weights were assigned to solutions based on a multivariate Gaussian likelihood, with means, standard deviations, and covariance taken from Table 1. In the left panel 1(a), an orange distribution is shown in addition to the blue distribution. Both blue distributions in 1(a) and 1(b) assume $\mathrm{cov}(J_{2},J_{4})=0$, whereas the orange distribution for Uranus takes $\mathrm{cov}(J_{2},J_{4})=0.9861\,\sigma_{J_{2}}\sigma_{J_{4}}$.
@@
Figure 2: Solution space distribution (2(a)) and contour of the distribution (2(b)) for Uranus. The left panel 2(a) shows the weighted distribution of all successful uncorrelated Uranus density profiles. The inset shows a zoomed in view of the outermost region, from $r/R=1$ to $r/R=0.7$. Note that the zoom is not aspect-preserving. The colour scale shows the relative frequency over all distributions of a certain density value at a given radius. Coloured lines show density profiles from previous studies for comparison. Solid and dashed lines correspond to models U1 & U2 from Nettelmann et al. (2013) (orange), U1 & U3 from Morf & Helled (2025) (green), and V2 & V3 from Vazan & Helled (2020) (blue), respectively. The colour scale on the right panel 2(b) shows the percentile intervals in 10% increments, except the last interval, where a 95% interval was added. The black line shows the median. The dashed lines show the 16th (lower) and 84th (upper) percentiles.
@@
Figure 3: Same as Figure 2 but for Neptune. The solid and dashed line in the left panel 3(a) correspond to the models N1 and N2b from Nettelmann et al. (2013) (orange), and to models N1 and N3 from Morf & Helled (2025) (green), respectively.
@@ 3 Results
Below we present the inferred density profiles and their characteristics. For simplicity, we refer to the normalised volumetric mean radius of a spheroid simply as ‘the normalised radius $r/R$’.
@@ 3.1 Density profiles
First, we present an aggregate view of the inferred profiles from the successful runs. A run is considered to be successful if its resulting density profile did not exceed $\rho_{max}$ and achieved the observed gravitational moments within the multivariate three standard deviation threshold. All runs within this threshold were then weighted according to their multivariate Gaussian likelihood to recover a distribution compatible with the data of Table 1. Figure 1 shows the gravitational moment distributions and Figures 2 and 3 show the density profiles as a distribution in logarithmic scale and contour respectively. Figure 4 shows their central densities while Figure 5 shows the size of the confidence intervals (68% and 95%).
@@
The gravitational moment distributions shown in Figure 1 demonstrate that our results cover the whole $J_{2}-J_{4}$ space faithfully, especially so for Uranus. For Neptune, the uncertainties are so large in comparison, that our algorithm struggled to find a significant amount of solutions outside the multivariate one standard deviation threshold, especially so for $J_{4}$. Since French et al. (2024) also provided a covariance estimate in their results, we additionally show a correlated distribution (orange) in panel 1(a) for Uranus. While the change between the uncorrelated and correlated distribution looks significant in this panel, the impact on subsequent results is negligible. All subsequent plots hence exclusively display results for the uncorrelated distributions to enable a fair comparison between Uranus and Neptune. Correlated Uranus results produce visually indistinguishable plots.
@@
Figures 2 and 3 show the density solution space for Uranus and Neptune, respectively. We find that successful density profiles are least constrained in the planetary deep interior. The outermost regions are highly constrained, with many distributions being affected by the minimum density increase. Around $\sim$85% of the planetary radius ($r/R=0.85$), a more diverse regime takes over for both planets. Still, the general shape continues to show a relatively tight constriction of the possible space up to a normalised radius of $\sim 0.4R$. This is most apparent in Figure 5. Notably, there is a bump at around $r/R=0.65$ for Uranus (Figure 2) and $r/R=0.7$ for Neptune (Figure 3). In the innermost region, a large increase in the space of possible distributions develops. The logarithmic scale is somewhat misleading, as can be seen by consulting the contour plots 2(b) and 3(b). Though many distributions fall outside the 95% region (especially above it), their relative frequency is small, on the order of $10^{-4}$ or less. The median density (as a function of radius) for Uranus and Neptune shows gradual increases in density, consistent with composition gradients. Note that this median density shown in panels 2(b) and 3(b) is neither an actual density profile from the solution ensemble nor a maximum-likelihood indicator.
@@
We next compare our results to previous work, first focusing on models in the literature that are based on equations of state and hence infer composition and temperature alongside density (and pressure). We find that the central densities of profiles from Helled et al. (2011) and Vazan & Helled (2020) are concentrated around the peak of the distribution, not its median. Profiles from Nettelmann et al. (2013) have core densities higher than the 84th percentile. More recent work by Militzer (2024) is similar, predicting core densities that are close to those inferred by Nettelmann et al. (2013). The Uranus models U1 and U3 with extensive composition gradients from Morf & Helled (2025) are close to the median result. Their convective models U2 and U4 are more similar to Vazan & Helled (2020) and are situated at the peak of the distribution. We note that although their U2 and U4 profiles possess the same comparatively low central density of $\sim 4.5$ g/cm3, their inferred total rock-to-water mass fractions vary significantly (0.64 vs. 3.92). We can therefore conclude that the density-pressure relation alone is insufficient for determining the planetary bulk composition. Interestingly, the N1 profile for Neptune by Morf & Helled (2025) falls outside the 95th percentile with a significantly higher central density. This is due to their approach that can also probe less-standard solutions. On the other hand, their N3 solution is close to the median result.
@@
Studies that, like this work, analyse large ensembles of interior profiles enable a more direct and systematic comparison. For example, Figure 4 in Helled & Fortney (2020), Figure 10 in Movshovitz & Fortney (2022), and Figure 2 in Neuenschwander & Helled (2022) are directly comparable with Figures 2(b) and 3(b). Helled & Fortney (2020) employed eighth-order polynomials, and Movshovitz & Fortney (2022) additionally incorporated sigmoid functions. They obtained profiles with generally higher densities in the deep interior for both Uranus and Neptune, especially in the case of Helled & Fortney (2020) below $0.3R$. Their median profiles also flatten off below $0.1R$, whereas our results show no such trend. Our medians maintain a consistent slope below $0.3R$ and even curve up slightly. Moreover, the distribution of Helled & Fortney (2020) is symmetric around the median, whereas our results show a wider spread for higher-density solutions. The inclusion of sigmoid terms in Movshovitz & Fortney (2022) produced extended flat regions that do not appear in our models due to the imposed minimum density increment per layer, $\Delta_{\min}$. In the work of Neuenschwander & Helled (2022), interior structures were constructed using three consecutive polytropes. This imprinted distinct features on the confidence intervals, including sharp discontinuities and narrow choke points, which are absent in our results. Similar to our results, their approach yielded a distribution with a wider spread around the median for higher central densities.
The wider spread for higher central densities is further illustrated in Figure 4. We find that the central densities are very unequally distributed, rising to a peak rapidly and gradually descending as density increases for both planets. For Uranus, the peak is very tight at $\rho\sim$ 4500 kg m-3, but the median is still somewhat higher. The solutions for Neptune are distributed more broadly. Because of this, a significant portion of solutions exceeded the imposed maximum density $\rho_{max}$ and were excluded. The peak of the distribution is at $\rho\sim$ 5200 kg m-3.
@@
Figure 5 reflects the span of our results. The uncertainty of the gravitational moment measurements of Neptune is a factor of ten larger than that of Uranus (Table 1). Once could therefore expect that our results are considerably less constrained for Neptune as well. However, Figure 5 shows that this is not the case. Our density profile results are only slightly more widely spread for Neptune than they are for Uranus. Tighter bounds in $J_{2}$-$J_{4}$ space do not linearly translate to tighter bounds in the density solution space. Even if the $J_{2}$ and $J_{4}$ values were perfectly known, the density profile solution space would remain degenerate. Only a measurement of all the gravitational moments $\left\{J_{2n}\mid n\in\mathbb{N}\right\}$ would provide a unique solution for a north-south and azimuthally symmetric planet.
Finally, note that in the contour (Figures 2(b) and 3(b)) and confidence interval (Figure 5) plots, small ‘spikes’ are visible concentrated around inverses of multiples of 2, most notably at $r/R=0.5$. These are artefacts of the initial density profile generation method and are addressed in Section 4 as a caveat.

### v3

@@
We present an algorithm to efficiently sample the full space of planetary interior density profiles. Our approach uses as few assumptions as possible to pursue an agnostic algorithm. The algorithm avoids the common Markov chain Monte Carlo method and instead uses an optimisation-based gradient-descent approach designed for computational efficiency. In this work, we use Uranus and Neptune as test cases and obtain empirical models that provide density and pressure profiles consistent with the observed physical properties (total mass, radius, and gravitational moments). We compared our findings to other work and find that while other studies are generally in line with our findings, they do not cover the entire space of solutions faithfully. Furthermore, we present guidance for modellers that construct Uranus or Neptune interior models with a fixed number of layers. We provide a statistical relation between the steepness classifying a density discontinuity and the resulting number of discontinuities to be expected. For example, if one classifies a discontinuity as a density gradient larger than 0.02 kg m-4, then most solutions should have at most one such discontinuity. Finally, we find that discontinuities, if present, are concentrated around a planetary normalised radius of 0.65 for Uranus and 0.7 for Neptune. Our algorithm to efficiently and faithfully investigate the full space of possible interior density profiles can be used to study all planetary objects with gravitational field data.
@@ 1 Introduction
Uranus and Neptune have always been key to understanding the formation and evolution history of the Solar System. As the investigation of our own planetary system continues to expand, the study of exoplanets is gaining more attention. Although the characterisation of exoplanets remains limited, it is now clear that in terms of mass and size, many exoplanets in the galaxy resemble Uranus and Neptune (Zhu & Dong, 2021). By revealing the internal structures of Uranus and Neptune, we can improve our knowledge of these two planetary objects in our own neighbourhood and the many observed intermediate-mass exoplanets orbiting distant stars.
To study the internal structure of a planet, we need to know its density profile. In the absence of direct measurements, the profile can be inferred from other observations. Density profiles can be inferred by varying all values constituting a profile until the result matches the available observational data (for example total mass, radius, and the gravitational field). The number of such observational parameters is typically small; for Uranus and Neptune, there are fewer than ten. In contrast, density profiles typically consist of orders of magnitude more data points. Inferring density profiles is therefore severely degenerate, as many density profiles can fit the same observational data.
To address this problem, one can simplify the approach used to infer density profiles by relying on a multitude of assumptions. However, caution must be taken. Making unsound simplifying assumptions could exclude a large portion (or all) of the valid density profiles. Therefore, obtaining a complete overview of all possible solutions is paramount before simplifications are even considered.
@@
The current numerical method of choice for empirical structure modelling is often the Markov chain Monte Carlo (MCMC) method (for example Marley et al., 1995; Movshovitz & Fortney, 2022). The MCMC approach starts by generating initial density profiles and then randomly alters their parameters until they fit the observational data. This poses challenges due to the aforementioned large number of data points that make up a density profile. Therefore, one must either tolerate the slow convergence of the MCMC method when handling countless parameters or reduce the dimensionality of the parameter space to accelerate the search. Previous studies have taken the latter route by parametrising density profiles with polytropes, polynomials, or sigmoid functions. This reduces the number of parameters to an order of magnitude comparable to the number of observational parameters, but it can introduce biases, which is precisely what we aim to avoid here.
We replace the common MCMC method with a new optimisation-based approach that uses an algorithm to efficiently investigate the full space of possible interior density profiles that fit the known observational data of a planet. The assumptions made for the profiles are reduced to an absolute minimum: We only require that there are no density plateaus or inversions and that the maximal density does not exceed the density of any feasible material. Otherwise, (with a resolution of $N=1024$) the $N$ density values are all allowed to change freely. Because we treat all $N$ density values as free parameters, the method avoids biases introduced by the parametrisations of previous MCMC studies. However, this alone does not ensure that the results are free from bias. We must further ensure that the generation of starting profiles is unbiased, and — because we no longer use an MCMC algorithm — we must additionally ensure that the way we find the density profiles is unbiased. For the first point, if the generation of starting profiles favours certain types of density profiles, for instance, by initially under-representing profiles with sharp density discontinuities, an adjustment that increases the sampling of such profiles is required. For the second point, we need to verify that our optimisation approach does not overlook harder-to-find local solutions in favour of dominant global ones that are easily spotted.
Our paper is structured as follows. Section 2 provides an overview of the algorithm. In Section 3 we present and interpret the results. In Section 4 we discuss the results, highlight caveats, and provide an outlook for future work. Section 5 presents a summary of the key findings of this study. We note that our methodology is based on the theory of figures (ToF, Zharkov & Trubitsyn, 1978). For the purposes of notation, we follow the elaboration presented in Appendix A of Morf et al. (2024). We employed the open-source code PyToF [^1] See https://doi.org/10.5281/zenodo.17592269 (Morf, 2025) to implement the ToF numerically.
@@ 2.1 Algorithm
Planets are flattened due to their rotation. Surfaces of both constant density and potential are not spherical shells defined by a single radius, $r$, but are rather spheroidal in shape. Spheroidal surfaces are described by a function, $r_{l}(\vartheta)$, that depends on the polar angle, $\vartheta$. The variable $l$ uniquely identifies a spheroid and represents its volumetric mean radius:
@@
where $R_{\text{ref}}$ is a normalisation constant, $m$ denotes the planetary mass, and $P_{2n}$ are the Legendre polynomials. The $J_{2n}$ are called gravitational moments and can be calculated with the ToF.
For the numerical implementation, density profiles consist of $N$ density values, $\rho_{k}$, at evenly spaced level surfaces, $l_{k}$. We restricted the maximum density to some threshold, $\rho_{max}$, and enforced a minimum density increase per step, $\Delta_{\text{min}}$. The density values are derived from N1N-1 parameters, $p_{i}$, as
@@
The weights, $w_{i}$, were added to normalise the parameter values and are defined as
@@
The scaling factor $\alpha$ was added to ensure the density profile mass, $m_{calc}$, agrees with the planetary mass, $m$. It is defined as
@@
where the calculated mass, $m_{calc}$, was obtained by numerically integrating the density values over the level surfaces via
@@
Listing 1: The optimisation algorithm.
@@
Given the number of epochs to run the algorithm for, we first created some suitable starting parameters. Then, for each epoch, we checked whether the optimisation converged to a stable solution. If not, we performed epochsize for many optimisation steps. For each optimisation step, we calculated the gradient and moved along it according to the Adam gradient descent scheme (Kingma & Ba, 2017). Then, we projected the new parameters to enforce our minimum density increase. Once convergence was detected, the final result and other data, was returned. Further details on the algorithm can be found in the Appendix A. Explanations regarding the requirements of an unbiased generation method and optimisation method are given in Appendix C. The method’s limitations are discussed in Section 4.
Table 1: Physical values for Uranus and Neptune.
@@ 2.2 Simulation setup
The simulation was performed on an AMD EPYC 7702 Hypervisor CPU with 96 cores over 120 hours. We dedicated 64 cores to Uranus and 32 towards Neptune. We used twice as many cores for Uranus in order to run correlated ($\text{cov}(J_{2},J_{4})=0.9861\sigma_{J_{2}}\sigma_{J_{4}}$; see Table 1) and uncorrelated ($\text{cov}(J_{2},J_{4})=0$) simulations in parallel. All physical values for Uranus and Neptune are given in Table 1. The number of layers, $N$, was set to 1024, and the epoch size was 25 steps. We chose a maximum of 120 epochs (3000 steps). The scalar factor, $c$, introduced in Equation 10 was set to $100$. This factor ensures that the optimisation cost function does not consider extreme values for $\alpha$ (far away from $\alpha=1$) while optimising for gravitational field data. As a convergence criterion within the ToF algorithm PyToF, we employed $\left|J_{2n}^{i+1}-J_{2n}^{i}\right|\leq 10^{-10}J_{2n}^{i}$ to check if convergence was achieved after the $i$-th iteration. The maximum acceptable density $\rho_{max}$ was set to 20000 kg m-3. The minimum acceptable density increase per unit length, $\Delta_{\text{min}}$, was set to 1 $\Delta\ell^{-1}$ kg m-3, where $\Delta\ell=R/N$. Explicitly, this resulted in a value of $\Delta\ell=$ 24960 m for Uranus and $\Delta\ell=$ 24186 m for Neptune (rounded to the nearest meter). This yielded $\Delta_{\text{min}}=$ 4.01 $\cdot$ 10-5 kg m-4 for Uranus and $\Delta_{\text{min}}=$ 4.13 $\cdot$ 10-5 kg m-4 for Neptune (rounded to two significant figures).
@@
Overall, 694465 (791901) runs were completed for the correlated (uncorrelated) Uranus data, and 731407 runs were completed for Neptune. For Uranus, 90.39% (90.48%) of the runs respected $\rho_{max}$, and 18.77% (18.56%) achieved gravitational moments, $J_{2n}$, within the multivariate three-standard-deviation threshold, which is 20.77% (20.56%) of those that respected $\rho_{max}$. For Neptune, the numbers (in the same order) are 89.22% and 88.83%, which is 99.57% of those that respected $\rho_{max}$. Our optimisation-based approach and the developed algorithm are highly efficient: The average time to complete a run was approximately 20 seconds.
@@
We note that throughout all subsequent results, the parameter space of the possible density profiles for Neptune is larger in comparison to Uranus. This is because the measurements we have for the gravitational moments of Uranus are about one order of magnitude more precise than those for Neptune (Table 1). Consequently, Uranian runs had a $\sim$19% success rate, compared to $\sim$89% of the Neptunian runs.
@@
Figure 1: Gravitational moment distribution for successful Uranus (1(a)) and Neptune (1(b)) profiles. Both panels include a 2D histogram where bins more frequently occupied by solutions are shown with a relatively darker colour. Additionally, 1D histograms display the weighted number of solutions along each axis. The weights were assigned to solutions based on a multivariate Gaussian likelihood, with means, standard deviations, and covariance taken from Table 1. In the left panel 1(a), an orange distribution is shown in addition to the blue distribution. Both blue distributions in 1(a) and 1(b) assume $\mathrm{cov}(J_{2},J_{4})=0$, whereas the orange distribution for Uranus takes $\mathrm{cov}(J_{2},J_{4})=0.9861\,\sigma_{J_{2}}\sigma_{J_{4}}$.
@@
Figure 2: Solution space distribution (2(a)) and contour of the distribution (2(b)) for Uranus. The left panel 2(a) shows the weighted distribution of all successful uncorrelated Uranus density profiles. The inset shows a zoomed-in view of the outermost region, from $r/R=1$ to $r/R=0.7$. Note that the zoom is not aspect-preserving. The colour scale shows the relative frequency over all distributions of a certain density value at a given radius. Coloured lines show density profiles from previous studies for comparison. Solid and dashed lines correspond to models U1 and U2 from Nettelmann et al. (2013) (orange); U1 and U3 from Morf & Helled (2025) (green); and V2 and V3 from Vazan & Helled (2020) (blue), respectively. The colour scale in the right panel 2(b) shows the percentile intervals in 10% increments, except the last interval, where a 95% interval was added. The black line shows the median. The dashed lines show the 16th (lower) and 84th (upper) percentiles.
@@
Figure 3: Same as Figure 2 but for Neptune. The solid and dashed line in the left panel 3(a) correspond to the models N1 and N2b from Nettelmann et al. (2013) (orange) and to models N1 and N3 from Morf & Helled (2025) (green), respectively.
@@ 3 Results
Below we present the inferred density profiles and their characteristics. For simplicity, we refer to the normalised volumetric mean radius of a spheroid as ‘the normalised radius $r/R$’.
@@ 3.1 Density profiles
First, we present an aggregate view of the inferred profiles from the successful runs. A run is considered to be successful if its resulting density profile did not exceed $\rho_{max}$, and it achieved the observed gravitational moments within the multivariate three standard deviation threshold. All runs within this threshold were then weighted according to their multivariate Gaussian likelihood to recover a distribution compatible with the data of Table 1. Figure 1 shows the gravitational moment distributions, and Figures 2 and 3 respectively show the density profiles as a distribution in logarithmic scale and contour. Figure 4 shows their central densities, while Figure 5 shows the size of the confidence intervals (68% and 95%).
@@
The gravitational moment distributions shown in Figure 1 demonstrate that our results cover the whole $J_{2}-J_{4}$ space faithfully, especially for Uranus. For Neptune, the uncertainties are so large in comparison that our algorithm struggled to find a significant amount of solutions outside the multivariate threshold of one standard deviation, especially for $J_{4}$. Since French et al. (2024) also provide a covariance estimate in their results, we additionally show a correlated distribution (orange) in panel 1(a) for Uranus. While the change between the uncorrelated and correlated distribution looks significant in this panel, the impact on the subsequent results is negligible. All subsequent plots hence exclusively display results for the uncorrelated distributions to enable a fair comparison between Uranus and Neptune. The correlated Uranus results produce visually indistinguishable plots.
@@
Figures 2 and 3 show the density solution space for Uranus and Neptune, respectively. We find that successful density profiles are least constrained in the planetary deep interior. The outermost regions are highly constrained, with many distributions being affected by the minimum density increase. Around $\sim$85% of the planetary radius ($r/R=0.85$), a more diverse regime takes over for both planets. Still, the general shape continues to show a relatively tight constriction of the possible space up to a normalised radius of $\sim 0.4R$. This is most apparent in Figure 5. Notably, there is a bump at around $r/R=0.65$ for Uranus (Figure 2) and $r/R=0.7$ for Neptune (Figure 3). In the innermost region, a large increase in the space of possible distributions develops. The logarithmic scale is somewhat misleading, as can be seen by consulting the contour plots 2(b) and 3(b). Though many distributions fall outside the 95% region (especially above it), their relative frequency is small, on the order of $10^{-4}$ or less. The median density (as a function of radius) for Uranus and Neptune shows gradual increases in density, consistent with composition gradients. We note that the median density shown in panels 2(b) and 3(b) is neither an actual density profile from the solution ensemble nor a maximum-likelihood indicator.
@@
Next we compare our results to previous work, focusing first on models in the literature that are based on equations of state and hence infer composition and temperature alongside density (and pressure). We find that the central densities of profiles from Helled et al. (2011) and Vazan & Helled (2020) are concentrated around the peak of the distribution, not its median. Profiles from Nettelmann et al. (2013) have core densities higher than the 84th percentile. The more recent work by Militzer (2024) is similar, predicting core densities that are close to those inferred by Nettelmann et al. (2013). The Uranus models U1 and U3 with extensive composition gradients from Morf & Helled (2025) are close to the median result. Their convective models U2 and U4 are more similar to Vazan & Helled (2020) and are situated at the peak of the distribution. We note that although their U2 and U4 profiles possess the same comparatively low central density of $\sim 4.5$ g/cm3, their inferred total rock-to-water mass fractions vary significantly (0.64 vs. 3.92). We can therefore conclude that the density-pressure relation alone is insufficient for determining the planetary bulk composition. Interestingly, the N1 profile for Neptune by Morf & Helled (2025) falls outside the 95th percentile with a significantly higher central density. This occurs because their approach can also probe less standard solutions. On the other hand, their N3 solution is close to the median result.
Studies that, as in this work, analyse large ensembles of interior profiles enable a more direct and systematic comparison. For example, Figure 4 in Helled & Fortney (2020), Figure 10 in Movshovitz & Fortney (2022), and Figure 2 in Neuenschwander & Helled (2022) are directly comparable to Figures 2(b) and 3(b). Helled & Fortney (2020) employed eighth-order polynomials, and Movshovitz & Fortney (2022) additionally incorporated sigmoid functions. They obtained profiles with generally higher densities in the deep interior for both Uranus and Neptune, especially in the case of Helled & Fortney (2020), below $0.3R$. Their median profiles also flatten off below $0.1R$, whereas our results show no such trend. Our medians maintain a consistent slope below $0.3R$ and even curve up slightly. Moreover, the distribution of Helled & Fortney (2020) is symmetric around the median, whereas our results show a wider spread for higher-density solutions. The inclusion of sigmoid terms in Movshovitz & Fortney (2022) produced extended flat regions that do not appear in our models due to the imposed minimum density increment per layer, $\Delta_{\min}$. In the work of Neuenschwander & Helled (2022), interior structures were constructed using three consecutive polytropes. This imprinted distinct features on the confidence intervals, including sharp discontinuities and narrow choke points, which are absent in our results. Similar to our results, their approach yielded a distribution with a wider spread around the median for higher central densities.
@@
The wider spread for higher central densities is further illustrated in Figure 4. We find that the central densities are very unequally distributed, rising to a peak rapidly and gradually descending as density increases for both planets. For Uranus, the peak is very tight at $\rho\sim$ 4500 kg m-3, but the median is still somewhat higher. The solutions for Neptune are distributed more broadly. Hence, a significant portion of solutions exceeded the imposed maximum density $\rho_{max}$ and were excluded. The peak of the distribution is at $\rho\sim$ 5200 kg m-3.
@@
Figure 5 reflects the span of our results. The uncertainty of the gravitational moment measurements of Neptune is a factor of ten larger than that of Uranus (Table 1). One could therefore expect that our results are considerably less constrained for Neptune as well. However, Figure 5 shows that this is not the case. Our density profile results are only slightly more widely spread for Neptune than they are for Uranus. Tighter bounds in the $J_{2}$-$J_{4}$ space do not linearly translate to tighter bounds in the density solution space. Even if the $J_{2}$ and $J_{4}$ values were perfectly known, the density profile solution space would remain degenerate. Only a measurement of all the gravitational moments $\left\{J_{2n}\mid n\in\mathbb{N}\right\}$ would provide a unique solution for a north-south and azimuthally symmetric planet.
Finally, we note that in the contour (Figures 2(b) and 3(b)) and confidence interval (Figure 5) plots, small ‘spikes’ are visible concentrated around inverses of multiples of two, most notably at $r/R=0.5$. These are artefacts of the initial density profile generation method and are addressed in Section 4 as a caveat.